Suppose a test given to students has a mean of 500 and standard deviation of 50 and the test scores are normally distributed.
What is the probablity that a single random student has a score between 495 and 503?
What is the shape, mean and standard deviation of the sampling distribution of mean(x bar) is based of a sample of n=36 randomly selected students?

Answer by stanbon(75887)

Suppose a test given to students has a mean of 500 and standard deviation of 50 and the test scores are normally distributed.
What is the probablity that a single random student
has a score between 495 and 503?
--------
z(495) = (495-500)/50 = -1/10
z(503) = (503-500)/50 = 3/50
--------
P(495< x < 503) = P(-0.1 < z < 0.06) = 0.0638
====================
Cheers,
Stan H.
===================


What is the shape, mean and standard deviation of the sampling distribution of mean(x bar) is based of a sample of n=36 randomly selected students?